OpenStudy (anonymous):

Solve the equation (-5/6)x = 40

pooja195 (pooja195):

Multiply by the reciprocal \[\huge~\rm~\bf~\frac{ -5 }{ 6 }x~\times \frac{ 6 }{ -5 }x\]

pooja195 (pooja195):

You are left with x= ? \[\huge~\rm~\frac{ 6 }{ -5 }\times 40=? \]

OpenStudy (anonymous):

= -4.8?

OpenStudy (anonymous):

-48

pooja195 (pooja195):

Correct :)